OpenStudy (anonymous):

MEDAL Which of the following describes the function −x3 + 5?

OpenStudy (anonymous):

A. The degree of the function is odd, so the ends of the graph continue in opposite directions. Because the leading coefficient is positive, the left side of the graph continues down the coordinate plane and the right side continues upward. B. The degree of the function is odd, so the ends of the graph continue in the same direction. Because the leading coefficient is negative, the left side of the graph continues down the coordinate plane and the right side also continues downward. C. The degree of the function is odd, so the ends of the graph continue in opposite directions. Because the leading coefficient is negative, the left side of the graph continues up the coordinate plane and the right side continues downward. D. The degree of the function is odd, so the ends of the graph continue in the same direction. Because the leading coefficient is positive, the left side of the graph continues up the coordinate plane and the right side also continues upward.

OpenStudy (mathmale):

I think you mean\[-x^3 + 5\] or \[f(x)=5-x^3\]

OpenStudy (mathmale):

Please read thru the several answer choices. Weed out those that make the least sense to you. Come up with reasons for selecting the choice you think is most likely to be correct.

OpenStudy (mathmale):

In your shoes I would graph y=x^3 first. Next, I'd change the sign to get z=-x^3. Next, I'd form the new function h(x) by adding 5 to z: \[z=5-x^3\] Each time I made a change, I'd ask myself how that change affected the graph.

OpenStudy (anonymous):

Yes, I meant x\[-x^{3}\]

OpenStudy (anonymous):

OKAY OPEN STUDY DOESN'T WORK IT KEEPS LOADING FOR LIKE 800 HOURS I GUESS IM JUST GOING TO FAIL. LOL.